OpenStudy (anonymous):

√3 (√6 + √3) + (√8 - 5)

OpenStudy (ash2326):

\[\sqrt(3)(\sqrt(6)+\sqrt(3))+(\sqrt(8)-5)\] \[\sqrt(18)+3+\sqrt(8)-5\] \[-2+3\sqrt2+2\sqrt2\] \[-2+5\sqrt2\]